# The Class 0 protostars in Orion: Characterizing the properties of their magnetized envelopes

### v1

@@ 3.4 Fragmentation Level and Disk Properties
Figure 4: The stacked histogram of disk parameters for protostars with different B-field structure. Upper row: disk mass $M_{\rm disk}$ (panel A), disk radius $R_{\rm disk}$ (B). Bottom row: disk inclination angle $\theta_{\rm Inc}$ (panel C and D). Dashed and solid lines indicate the mean and median values of the distribution. In panel D, “Face-on”, “Inclined”, and “Edge-on” indicate the inclination angle within a range of 0–30, 30–60, and 60–90, respectively. The disk mass, disk radius, and disk inclination angle are obtained from Tobin et al. (2020).
@@
Tobin et al. (2020) identified protostellar disks in the Orion molecular clouds using both VLA and ALMA observations, with further refinements presented in their subsequent study of Tobin et al. (2022). All objects in our sample are included within the VANDAM survey, which enables us to estimate the multiplicity of these systems. Palau et al. (2013) defined the fragmentation level as the number of millimeter sources inside a region with a radius of 0.05 pc for massive dense cores, whereas Busquet et al. (2016) used a different radius of 0.15 pc. We follow the above methods, but define a different radius of 1200 au, which is the scale for estimating the physical parameters above. Then the number of fragments (including small-scale fragmentation, that was only resolved in the survey with higher resolution of $\sim$40–400 au-scale companions) identified from the VANDAM survey within this innermost envelope region is indicative of the fragmentation level $N_{\rm F}$. It is possible that the outer ($>1200$ au) regions may influence the fragmentation of the envelopes, given that the typical scale for the star-forming envelopes is up to several thousand au (e.g., Myers et al., 2000). We thus define larger radii of 1800 au and 2400 au for estimating extra contribution of $N_{\rm F}$, but with less weighting values of 0.5 and 0.25, respectively. As shown in Figure 3, which shows the 0.87 mm dust emission map of an example HOPS-78, a protostar (black asterisk) is located in the innermost region with a radius of 1200 au (black dashed circle), then $N_{\rm F}$ in this region is counted as 1. While there are 2 (blue asterisks) and 1 (red asterisk) protostars in annular regions with inner and outer radii of 1200 and 1800 au (blue dashed circle), and 1800 and 2400 au (red dashed circle), then the $N_{\rm F}$ for each region is counted as 1 and 0.25, respectively. The total $N_{\rm F}$ for HOPS-78 is therefore assigned a value of 2.25. Column 11 of Table LABEL:Tab:parameters lists the fragmentation level for the sample.
@@
The mass, radius, and inclination angle of the disks listed in column 12–14 of Table LABEL:Tab:parameters are also taken from Tobin et al. (2020). The upper row of Figure 4 shows the distribution of disk mass and radius, with median and mean values of 118.5 and 171.7 M, and 69.7 and 79.2 au, respectively. While the bottom row of Figure 4 presents the distribution of disk inclination angle, which is approximated as $\theta_{\rm inc}=\arccos(\Theta_{\rm min}/\Theta_{\rm maj})$, where $\Theta_{\rm maj}$ and $\Theta_{\rm min}$ are the deconvolved major and minor axes of the disk. Here the face-on and edge-on configurations are defined with an inclination angle of $\theta_{\rm Inc}$ = 0, and $\theta_{\rm Inc}$ = 90, respectively. The mean and median values of the inclination angle are 44.3 and 45.9, respectively. Note that the finite resolution and vertical thickness of the disks (Sheehan et al., 2022; Lin et al., 2023; van’t Hoff et al., 2023) will bias this simple inclination estimate against 90, which is probably why it peaks at $\sim 45^{\circ}$ and not 60.
@@ 4 Discussion
In Section 3, we have derived the physical parameters of dust temperature, envelope gas mass, gas column density, and gas number density for the entire sample of BOPS within the inner region of 6″ (corresponding to a scale of 2400 au, with a radius of 1200 au). The distributions of these parameters are shown in Figure 2. We find from Figure 2(A–B) that protostars with high bolometric luminosity ($>$ 50 L) and/or envelope dust temperature ($>$ 25 K) tend to exhibit spiral or rotated-hourglass B-field morphology. In Figure 2(C–F), sources with compact polarized emission (self-scattering) tend to have relatively lower total flux density, envelope masses, gas column densities, and gas number densities, whereas protostars exhibiting a standard-hourglass B-field pattern tend to exhibit relatively larger values in these parameters. However, these standard-hourglass protostars have relatively low velocity dispersion, which is potentially indicative of the magnetic braking effect, whereby the strong field exerts a braking influence on the rotational rate (as shown in panel G of Figure 2). In Figure 2(H), we find standard-hourglass-B-field protostars have relatively low angle dispersion, in contrast to rotated-hourglass-, spiral-, and complex-field sources, which display a relatively larger angle dispersion.
@@
The disk properties are shown in Figure 4. We find from the first row that, the distribution of disk mass and radius appears random across different types of B-field morphologies. In the bottom row, in both types of sources dominated by self-scattering and exhibiting rotated-hourglass morphologies, with the mean polarization angle preferentially aligned parallel to the outflow direction, the disk inclination angle is found to be either very inclined or edge-on. However, in standard-hourglass sources, with the mean polarization angle oriented perpendicularly to the outflow, the disk inclination angle is typically found to be face-on. This is probably due to the fact that some disks in the VANDAM project are potentially influenced by the envelope emission (HOPS-87, HOPS-317S, and HOPS-403, see Tobin et al., 2020).
To statistically examine the relations between two pairs of parameters, we performed both the Pearson linear correlation coefficient ($r_{p}$) and Spearman’s rank correlation coefficient ($r_{s}$). The value of $r_{p}$ and that of $r_{s}$ are each bounded from -1 to 1, for negative to positive slope correlations, respectively. In each case, the closer the absolute value is to 1, the stronger the correlation. The p–value is used for the purpose of testing the null hypothesis that the two variables are uncorrelated. In general, statistically significant correlations are identified by p–value $<$ 0.05. Here we analyze the following figures, and based on these details, we will motivate a more detailed analysis on fragmentation and the possible effect of magnetic braking on disk formation.
@@ 4.1 Physical Envelope Properties
The dust temperature is expected to be correlated with the bolometric luminosity if the dust is primarily heated by the stellar radiation (see Section 3.1). During the early stages of star formation (Class 0 and early Class I), the protostar is deeply embedded within a dense envelope. In this phase, the luminosity is predominantly due to the accretion of material from the envelope through the disk onto the protostar, which releases gravitational energy. High-density regions are often associated with intense accretion activity as the protostar accumulates material. The inflow of material from the envelope onto the protostar leads to an increase in $L_{\rm bol}$. Additionally, a more massive envelope absorbs and re-radiates more energy and it is more efficient at cooling due to its higher optical depth. We plot the distribution between these physical parameters, as shown in Figure 5(A–F). However, our results reveal no significant correlations between $L_{\rm bol}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, and between $\overline{T}_{\rm env}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, as their correlation coefficients are close to 0, with p–values much higher than 0.05. The lack of a direct correlations are not surprising due to the dynamic nature of protostellar evolution. Protostellar envelope masses are likely influenced by both the final stellar mass of the protostar and the evolutionary stage, with envelope masses decreasing rapidly over time (Fischer et al., 2017). Luminosity, on the other hand, is influenced by episodic accretion, which may vary independently of the envelope mass (e.g., Zakri et al., 2022; Narang et al., 2024).
@@
Non-thermal motions, often driven by turbulence, are of significant importance in star-forming regions and affect the overall dynamics and evolution of the protostellar system. Therefore, $\sigma_{\rm nth}$ serves as an indicator of turbulence level, which enhance the accretion process by increasing the rate at which material is fed onto the protostar. As shown in Figure 5(G–H), a slight positive correlation is observed between $\sigma_{\rm nth}$ and $L_{\rm bol}$, as well as between $\overline{T}_{\rm env}$ and $L_{\rm bol}$. Protostars with higher value of $L_{\rm bol}$ typically drive powerful outflows and winds, injecting energy into the surrounding medium and thereby increasing the non-thermal velocity dispersion. In high turbulent regions, the non-thermal velocity dispersion provide support against gravitational collapse. Once, however, collapse is triggered, the turbulent motions channel material onto the forming protostar, influencing its luminosity. Theoretical models also suggest that higher levels of turbulence ($\sigma_{\rm nth}$) lead to more efficient accretion processes, leading to increased protostellar luminosities (e.g., Krumholz et al., 2005; Padoan et al., 2011; Federrath, 2015). Additionally, in regions with high $\sigma_{\rm nth}$, which are characterised by intense star formation activity (e.g., jets and winds), one might expect that higher dust temperatures may result from the combined effects of shock heating and enhanced radiative feedback from very young protostars. This is due to the fact that the surrounding dust is heated by protostellar outflows, winds and radiation.
@@
We note that these results will be influenced by the assumption of a typical temperature of 20 K. However, we find that 9 protostars exhibit relatively high values of $L_{\rm bol}$ and $\overline{T}_{\rm env}$. Of these, 3 display a rotated hourglass field shape (HOPS-168, HOPS-288, and HOPS-370), 5 of the protostars are spiral (HOPS-124, HOPS-182, HOPS-361N, HOPS-361S, and HOPS-384), and one protostar with compact polarized emission (HOPS-84).
@@ 4.2 Fragmentation
Figure 6(A–C) show the relation between $N_{\rm F}$ and $M_{\rm env}$, $\sigma_{\rm nth}$, and $\delta\phi$, respectively. It is important to note that the values of $N_{\rm F}$ are dominated by multiples within the inner 2400 au region, a scale that is consistent with the one used to estimate the physical parameters as discussed in Section 3. For protostars located in multiple systems with separation distances of less than 1200 au, the fragmentation level correlations are plotted only once, using the protostar with the highest dust emission as the reference point. Combined with the statistical test methods, we discuss the potential influence of specific factors on envelope fragmentation at the 2400 au scale in the following.
@@
Optical and near infrared studies suggest that the multiplicity rate drops with decreasing mass (e.g., Chini et al., 2012, 2013). We investigated the correlation between $N_{\rm F}$ and $M_{\rm env}$, and the result is shown in Figure 6(A). Based on the correlation coefficients shown in Figure 6(A), there is no clear correlation between $N_{\rm F}$ and $M_{\rm env}$. This result may suggest that the correlation observed in the optical/infrared is not yet established at the very early stages of star and cluster formation. It is possible, therefore, that the final number of members in a multiple system is determined after multiple star formation episodes, as suggested by both theoretical (e.g., Myers, 2011) and observational(e.g., Bik et al., 2012) studies. Furthermore, the process of fragmentation and the formation of multiple objects are typically governed by a complex interplay of factors beyond just the envelope mass. These interactions may lead to significant variability in outcomes, masking any simple correlation between envelope mass and multiplicity. Previous observational studies of massive dense cores suggest there is no correlation between mass and multiplicity within a radius of 0.05 pc of the dust peak (Palau et al., 2013).
@@
In recent decades, several studies on the turbulence scenario (e.g., Padoan et al., 2001; Vázquez-Semadeni et al., 2007) have demonstrated that turbulence is capable of generating density structures through the action of supersonic shocks, which compress and fragment the gas efficiently, particularly in the presence of significant turbulence. Subsequent studies have indicated that different types of turbulence, i.e. compressive and solenoidal, exert divergent effects on fragmentation (e.g., Federrath et al., 2008). Compressive turbulence is conducive to fragmentation, whereas solenoidal turbulence has the effect on slowing down fragmentation (e.g., Federrath et al., 2010; Girichidis et al., 2011). As shown in Figure 6(B), the fragmentation level $N_{\rm F}$ is weakly correlated with $\sigma_{\rm nth}$, and this correlation becomes more evident when the protostar HOPS-361S with $N_{\rm F}=4$ is excluded, with correlation coefficients of $r_{p}\sim 0.42$, $r_{s}\sim 0.40$, p–value $\ll 0.01$. This positive correlation between $N_{\rm F}$ and $\sigma_{\rm nth}$ in our result suggests that turbulence tends to promote the fragmentation, indicating that compressive turbulence may be more dominant on envelope scales. However, one should note that $\sigma_{\rm nth}$ may also include contributions from rotational and infalling motions on envelope scales, which cannot be distinguished with the limited spectral resolution of C17O. Therefore, deeper studies with sufficient spectral resolution should be carried out to reach more robust conclusions.
@@
On the other hand, the fragmentation level in massive dense cores could be explained by a B-field/turbulence balance (Palau et al., 2013) or changes in the mass-to-flux ratio (Busquet et al., 2016; Añez-López et al., 2020; Palau et al., 2021) on scales of several thousands au, suggesting a potential correlation between B-fields and fragmentation on envelope scales. One of the most longstanding methods for estimating the B-field strength is the Davis-Chandrasekhar-Fermi (DCF, Davis, 1999; Chandrasekhar et al., 1953) method, which relies on the principle that turbulent motions should visibly affect the distribution of the B-field in the plane of the sky. The basic form of the DCF formula is $B_{\rm{pos}}=\sqrt{4\pi\rho}\sigma_{\rm nth}/\delta\phi$, where $\rho$ is the gas density, and $\delta\phi$ is the measured angular dispersion in B-field position angles (as discussed in Section 3.3). Since $\sigma_{\rm nth}$ is barely resolved for most sources in our sample, which significantly affect the estimation of B-field strength. However, the above equation indicates that the field strength is inversely proportional to the dispersion in the polarization position angles, which provides a potential way to indirectly study the correlation between field strength and other physical parameters on envelope scales. In order to statistically test this correlation, we plot the distribution between the angle dispersion of the B-field $\delta\phi$ and the fragmentation level $N_{\rm F}$, as shown in Figure 6(C). Combing correlation coefficients, we find that $N_{\rm F}$ is positive correlated with $\delta\phi$. Even when excluding the outlier point of protostar HOPS-361S with $N_{\rm F}=4$, the positive correlation ($r_{p}\sim 0.50$, $r_{s}\sim 0.55$, p–value $\sim 0.03$) remains significant. This suggests that higher angle dispersion in the B-field is associated with higher multiplicity, and thus, the fragmentation tends to be suppressed by B-field. Previous studies have indicated that strong B-fields suppress fragmentation, as they constitute an additional form of support against gravitational contraction and also via the magnetic effect (inward pinching and gathering of the field lines dragged during collapse makes the magnetic strength larger at inner radius, and acts against rotation), which reduces the amount of angular momentum in the inner part of cores (e.g., Galli et al., 2006; Mellon et al., 2008; Hennebelle et al., 2008; Price et al., 2009; Myers et al., 2014; Cunningham et al., 2018; Kwon et al., 2019; Yen et al., 2021).
@@
In addition, multiplicity may also decrease with evolution from Class 0 to Class I (Tobin et al., 2020, 2022). However, since the BOPS sample predominantly consists of objects at the Class 0 stage, the sample size in our study is insufficient to differentiate evolutionary factors over time.

### v2

@@ 3.4 Fragmentation Level and Disk Properties
Figure 4: The stacked histogram of disk parameters for protostars with different B-field structure. Upper row: disk mass $M_{\rm disk}$ (panel A), disk radius $R_{\rm disk}$ (B). Bottom row: disk inclination angle $\theta_{\rm Inc}$ (panel C and D). Dashed lines in A, B and C panels indicate the mean values of the distribution for different B-field structures. In panel D, “Face-on”, “Inclined”, and “Edge-on” indicate the inclination angle within a range of 0–30, 30–60, and 60–90, respectively. The disk mass, disk radius, and disk inclination angle are obtained from Tobin et al. (2020).
@@
Tobin et al. (2020) identified protostellar disks in the Orion molecular clouds using both VLA and ALMA observations, with further refinements presented in their subsequent study of Tobin et al. (2022). All objects in our sample are included within the VANDAM survey, which enables us to estimate the multiplicity of these systems. Palau et al. (2013) defined the fragmentation level as the number of millimeter sources inside a region with a radius of 0.05 pc for massive dense cores, whereas Busquet et al. (2016) used a different radius of 0.15 pc. We follow the above methods, but define a different radius of 1200 au, which is the scale for estimating the physical parameters above. Then the number of fragments (including small-scale fragmentation, that was only resolved in the survey with higher resolution of $\sim$40–400 au-scale companions) identified from the VANDAM and the BOPS projects within the innermost envelope region is indicative of the fragmentation level $N_{\rm F}$. It is possible that the outer ($>1200$ au) regions may influence the fragmentation of the envelopes, given that the typical scale for the star-forming envelopes is up to several thousand au (e.g., Myers et al., 2000). We thus define larger radii of 1800 au and 2400 au for estimating extra contribution of $N_{\rm F}$, but with less weighting values of 0.5 and 0.25, respectively. For example, in Figure 3, which shows the 0.87 mm dust emission map of an example HOPS-78, a protostar (black asterisk) is located in the innermost region with a radius of 1200 au (black dashed circle), then $N_{\rm F}$ in this region is counted as 1. While there are 2 (blue asterisks) and 1 (red asterisk) protostars in annular regions with inner and outer radii of 1200 and 1800 au (blue dashed circle), and 1800 and 2400 au (red dashed circle), then the $N_{\rm F}$ for each region is counted as 1 and 0.25, respectively. Note that these protostars are identified from continuum maps of the BOPS and VANDAM survey. The total fragmentation level $N_{\rm F}$ for HOPS-78 is therefore assigned a value of 2.25. Column 11 of Table LABEL:Tab:paras lists the fragmentation level for 61 targets.
@@
The mass, radius, size, and inclination angle of the disks listed in column 5–8 of Table LABEL:Tab:parameters in the Appendix A are taken from Tobin et al. (2020). Figure 4 shows the distribution of 43 protostars. The rest 18 protostars lack information of the disk, or the disk dust emission is potentially contaminated by the envelope. The upper panels show the distribution of the disk mass and radius, while the bottom panel presents the distribution of disk inclination angle, which is estimated by $\theta_{\rm inc}=\arccos(\Theta_{\rm min}/\Theta_{\rm maj})$, where $\Theta_{\rm maj}$ and $\Theta_{\rm min}$ are the deconvolved major and minor axes of the disk. Here the face-on and edge-on configurations are defined with an inclination angle of $\theta_{\rm Inc}$ = 0–30, and $\theta_{\rm Inc}$ = 60–90, respectively. The trend of the distribution for disk properties, as well as other parameters, will be discussed in the following.
@@ 4 Discussion
In Section 3, we have derived physical parameters of envelope gas mass, gas column density, gas number density for entire sample of BOPS (61 protostars) based on the available bolometric luminosity (Furlan et al., 2017) and dust temperature (Lombardi et al., 2014), and non-thermal velocity dispersion for 60 protostars within the inner region of 6″ (corresponding to a scale of 2400 au, with a radius of 1200 au). Angle dispersion of B-field is also estimated within the same regions for 25 BOPS targets with sufficient Nyquist-sampling statistics. Combining with the VANDAM survey, the fragmentation level has been identified for 61 BOPS targets. In addition, disk properties have been presented for 43 BOPS targets from Tobin et al. (2020), without any contamination of envelope emission and other components of disk emission.
@@
The distributions of the these parameters are shown in Figure 2. Combining both the distribution and the mean values (dashed lines) for each type of B-field morphologies, we find that protostars with high bolometric luminosity ($>$ 50 L) and/or envelope dust temperature ($>$ 35 K) tend to exhibit spiral or rotated-hourglass B-field morphology (Figure 2(A–B)). In Figure 2(C–F), envelopes with compact polarized emission (self-scattering in blue) tend to have lower envelope masses, gas column densities, and gas number densities. Whereas protostars exhibiting a standard-hourglass B-field pattern tend to exhibit relatively larger values in these parameters. However, these standard-hourglass protostars have relatively low non-thermal velocity dispersion, which is potentially indicative of the magnetic braking effect, whereby the strong field exerts a braking influence on the rotational rate (as shown in panel G of Figure 2). In Figure 2(H), we find standard-hourglass-B-field protostars have relatively low angle dispersion, in contrast to rotated-hourglass-, spiral-, and complex-field sources, which display a relatively larger angle dispersion.
In Figure 4, which shows the distribution of disk properties, we find from the first row (panel A and B) that protostars exhibiting standard hourglass B-field structure have smaller disk size than those exhibiting other field configurations. In panel C, the mean values for protostars with different type of B-field structure are ranging from 25 to 50. Note that the finite resolution and vertical thickness of the disks (Sheehan et al., 2022; Lin et al., 2023; van’t Hoff et al., 2023) will bias this simple inclination estimate against 90, which would influence the location of the mean values. In panel D, in both types of sources dominated by self-scattering and exhibiting rotated-hourglass morphologies, with the mean polarization angle preferentially aligned parallel to the outflow direction, the disk inclination angle is found to be either very inclined or edge-on. However, in standard-hourglass sources, with the mean polarization angle oriented perpendicularly to the outflow, the disk inclination angle is typically found to be face-on.
Below we discuss the potential correlation between these parameters and disk properties. To statistically examine the relations between two pairs of parameters, we performed both the Pearson linear correlation coefficient ($r_{p}$) and the Spearman’s rank correlation coefficient ($r_{s}$). The values of $r_{p}$ and $r_{s}$ are each bounded from -1 to 1, for negative to positive slope correlations, respectively. In each case, the closer the absolute value is to 1, the stronger the correlation. The p–value is used for the purpose of testing the null hypothesis that the two variables are uncorrelated. In general, statistically significant correlations are identified by p–value $<$ 0.05. Here we analyze the following figures, and based on these details, we will motivate a more detailed analysis on fragmentation and the possible effect of magnetic braking on disk formation.
@@ 4.1 Physical Envelope Properties
The dust temperature is expected to be correlated with the bolometric luminosity if the dust is primarily heated by the stellar radiation (see Section 3.1). During the early stages of star formation (Class 0 and early Class I), the protostar is deeply embedded within a dense envelope. In this phase, the luminosity is predominantly due to the accretion of material from the envelope through the disk onto the protostar, which releases gravitational energy. High-density regions are often associated with intense accretion activity as the protostar accumulates material. The inflow of material from the envelope onto the protostar leads to an increase in $L_{\rm bol}$. Additionally, a more massive envelope absorbs and re-radiates more energy and it is more efficient at cooling due to its higher optical depth. 5(A–F) shows the distribution between these physical parameters for a sample of 55 envelopes (the remaining 6 targets have no bolometric luminosity information). However, our results reveal no significant correlations between $L_{\rm bol}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, and between $\overline{T}_{\rm env}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, as their correlation coefficients are close to 0, with p–values much higher than 0.05. The lack of direct correlations is likely because protostellar envelope masses are influenced by both the final stellar mass of the protostar and the evolutionary stage, with envelope masses decreasing rapidly over time (Fischer et al., 2017). Luminosity, on the other hand, is influenced by episodic accretion, which may vary independently of the envelope mass (e.g., Zakri et al., 2022; Narang et al., 2024).
@@
Non-thermal motions, often driven by turbulence, are of significant importance in star-forming regions and affect the overall dynamics and evolution of the protostellar system. Therefore, $\sigma_{\rm nth}$ serves as an indicator of turbulence level, which enhance the accretion process by increasing the rate at which material is fed onto the protostar. Protostars with higher value of $L_{\rm bol}$ typically drive powerful outflows and winds, injecting energy into the surrounding medium and thereby increasing the non-thermal velocity dispersion. In high turbulent regions, the non-thermal velocity dispersion provide support against gravitational collapse. Once, however, collapse is triggered, the turbulent motions channel material onto the forming protostar, influencing its luminosity. Theoretical models also suggest that higher levels of turbulence ($\sigma_{\rm nth}$) lead to more efficient accretion processes, leading to increased protostellar luminosities (e.g., Krumholz et al., 2005; Padoan et al., 2011; Federrath, 2015). Additionally, in regions with high $\sigma_{\rm nth}$, which are characterised by intense star formation activity (e.g., jets and winds), one might expect that higher dust temperatures may result from the combined effects of shock heating and enhanced radiative feedback from very young protostars. To test these senarios, we plot the distribution between $\sigma_{\rm nth}$ and $L_{\rm bol}$, and between $\sigma_{\rm nth}$ and $T_{\rm env}$ for 54 protostars (5 protostars lack luminosity information, while one protostar of HOPS-408 has no detection of velocity dispersion), as shown in Figure 5(G–H). A slight positive correlation is observed between $\sigma_{\rm nth}$ and $L_{\rm bol}$ from $r_{s}$ and p-value, as well as between $\overline{T}_{\rm env}$ and $L_{\rm bol}$ from $r_{p}$ and $r_{s}$. However, $r_{p}\sim 0$ in panel G and the p-value $>$ 0.05 in panel H suggest insignificant correlations between these parameters. The interplay between localized feedback, external turbulence, B-fields, and accretion variability makes it difficult to isolate a direct relationship, especially when dealing with deeply embedded young protostars in Orion, a highly turbulent environment. In addition, the limited spectral resolution would also affect these results. More focused studies on individual cores or higher-resolution observations might reveal more localized correlations that are otherwise masked by these factors.
@@
We also note that these results will be influenced by the assumption of a typical temperature of 20 K. However, we find that 7 protostars exhibit relatively high values $\overline{T}_{\rm env}$ ($>$ 40 K). Of these, 2 display a rotated hourglass field shape (HOPS-288 and HOPS-370), 5 of the protostars are spiral (HOPS-124, HOPS-182, HOPS-361N, HOPS-361S, and HOPS-384).
@@ 4.2 Fragmentation
Figure 6(A–C) show the relation between $N_{\rm F}$ and $M_{\rm env}$ for 61 protostars, between $N_{\rm F}$ and $\sigma_{\rm nth}$ for 60 protostars (HOPS-408 has no C17O detection, so its $\sigma_{\rm nth}$ cannot be resolved), and between $N_{\rm F}$ and $\delta\phi$ for 18 protostars out of 25 with enough polarzation statistics (the other 7 have a complex large-scale B-field morphology), respectively. It is important to note that the values of $N_{\rm F}$ are dominated by multiples within the inner 2400 au region, a scale that is consistent with the one used to estimate the physical parameters as discussed in Section 3. Combined with the statistical test methods, we discuss the potential influence of specific factors on envelope fragmentation at the 2400 au scale in the following.
@@
Optical and near infrared studies suggest that the multiplicity rate drops with decreasing mass (e.g., Chini et al., 2012, 2013), we thus investigated the correlation between $N_{\rm F}$ and $M_{\rm env}$. Based on the correlation coefficients shown in Figure 6(A), there is no clear correlation between $N_{\rm F}$ and $M_{\rm env}$. This result may suggest that the correlation observed in the optical/infrared is not yet established at the very early stages of star and cluster formation. It is possible, therefore, that the final number of members in a multiple system is determined after multiple star formation episodes, as suggested by both theoretical (e.g., Myers, 2011) and observational(e.g., Bik et al., 2012) studies. Furthermore, the process of fragmentation and the formation of multiple objects are typically governed by a complex interplay of factors beyond just the envelope mass. These interactions may lead to significant variability in outcomes, masking any simple correlation between envelope mass and multiplicity. Previous observational studies of massive dense cores also suggest that there is no correlation between mass and multiplicity within a radius of 0.05 pc of the dust peak (e.g., Palau et al., 2013).
@@
In recent decades, several studies on the turbulence scenario (e.g., Padoan et al., 2001; Vázquez-Semadeni et al., 2007) have demonstrated that turbulence is capable of generating density structures through the action of supersonic shocks, which compress and fragment the gas efficiently, particularly in the presence of significant turbulence. Subsequent studies have indicated that different types of turbulence, i.e. compressive and solenoidal, exert divergent effects on fragmentation (e.g., Federrath et al., 2008). Compressive turbulence is conducive to fragmentation, whereas solenoidal turbulence has the effect on slowing down fragmentation (e.g., Federrath et al., 2010; Girichidis et al., 2011). As shown in Figure 6(B), the fragmentation level $N_{\rm F}$ is uncorrelated with $\sigma_{\rm nth}$. However, it is important to note that $\sigma_{\rm nth}$ may include contributions from rotational and infalling motions on envelope scales, which cannot be distinguished with the limited spectral resolution of C17O. Therefore, deeper studies with sufficient spectral resolution should be carried out to reach more robust conclusions.
@@
On the other hand, the fragmentation level in massive dense cores could be explained by a B-field/turbulence balance (Palau et al., 2013) or changes in the mass-to-flux ratio (Busquet et al., 2016; Añez-López et al., 2020; Palau et al., 2021) on scales of several thousands au, suggesting a potential correlation between B-fields and fragmentation on envelope scales. One of the most longstanding methods for estimating the B-field strength is the Davis-Chandrasekhar-Fermi (DCF, Davis, 1999; Chandrasekhar et al., 1953) method, which relies on the principle that turbulent motions should visibly affect the distribution of the B-field in the plane of the sky. The basic form of the DCF formula is $B_{\rm{pos}}=\sqrt{4\pi\rho}\sigma_{\rm nth}/\delta\phi$, where $\rho$ is the gas density, and $\delta\phi$ is the measured angle dispersion of the B-field (as discussed in Section 3.3). Since $\sigma_{\rm nth}$ is barely resolved for most sources in our sample, which significantly affect the estimation of B-field strength. However, the above equation indicates that the field strength is inversely proportional to the dispersion in the polarization position angles, which provides a potential way to indirectly study the correlation between field strength and other physical parameters on envelope scales. In order to statistically test this correlation, we plot the distribution between the angle dispersion of the B-field $\delta\phi$ and the fragmentation level $N_{\rm F}$, as shown in Figure 6(C). Combing correlation coefficients, we find that $N_{\rm F}$ is positively correlated with $\delta\phi$. This suggests that higher angle dispersions in the B-field is associated with higher multiplicities, and thus envelope fragmentation tends to be suppressed by the B-field. Previous studies have indicated that strong B-fields suppress fragmentation, as they constitute an additional form of support against gravitational contraction and also via the magnetic effect (inward pinching and gathering of the field lines dragged during collapse makes the magnetic strength larger at inner radius, and acts against rotation), which reduces the amount of angular momentum in the inner part of cores (e.g., Galli et al., 2006; Mellon et al., 2008; Hennebelle et al., 2008; Price et al., 2009; Myers et al., 2014; Cunningham et al., 2018; Kwon et al., 2019; Yen et al., 2021).
@@
In addition, multiplicity may also decrease with the evolution from Class 0 to Class I (Tobin et al., 2020, 2022). However, since the BOPS sample predominantly consists of objects at the Class 0 stage, the sample size in our study is insufficient to differentiate evolutionary factors over time.